OpenStudy (anonymous):

How to get to the acceleration equation of an oscillating object, a=-w^2.x ? Thank you for your time kind people!

OpenStudy (vincent-lyon.fr):

Use Newton's second law of motion and Hooke's law for the restoring force.

OpenStudy (anonymous):

Ah yes thank you for the hint! F = ma and F = -kx, then I just rearrange. Ha I guess I just panicked.

OpenStudy (vincent-lyon.fr):

No problem :)

OpenStudy (anonymous):

I have seen it done as follows: x = A sin(w t) v = dx/dt = w A cos(w t) a = dv/dt = -(w^2) A sin(w t) = -(w^2) x

OpenStudy (anonymous):

That as well, yes :) Thank you very much!